A total of 285
tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was two times the number of adult tickets sold. How many adult tickets were sold?

Answer by ikleyn(52908)

.
Let x = # of adults tickets.


Then the number of student tickets is 2x.


x + 2x = 285


3x     = 285


x = 285/3 = 95.


ANSWER.  95 adult tickets.


CHECK.   95 + 2*95 = 285  tickets.   ! Correct !

Solved.